Let x be a continuous random variable that follows a normal distribution with a mean of 185 and a standard deviation of 18. Find the value of x so that the area under the normal curve between μ and x is approximately 0.4812 and the value of x is greater than μ.

1
Expert's answer
2021-12-06T11:24:18-0500
"P(Z<\dfrac{x-\mu}{\sigma})-0.5=0.4812"

"P(Z<\dfrac{x-\mu}{\sigma})=0.9812"

"\dfrac{x-\mu}{\sigma}\approx2.079189"

"x\approx\mu+2.079189\sigma"

Given "\mu=185, \sigma=18."


"x\approx185+2.079189(18)"

"x\approx222.425"